# Discovery of Critical Thresholds in Mixed Exposures and Estimation of Policy Intervention Effects using Targeted Learning

### v2

@@ Abstract
Traditional regulations of chemical exposure tend to focus on single exposures, overlooking the potential amplified toxicity due to multiple concurrent exposures. We are interested in understanding the average outcome if exposures were limited to fall under a multivariate threshold. Because threshold levels are often unknown a priori, we provide an algorithm that finds exposure threshold levels where the expected outcome is maximized or minimized. Because both identifying thresholds and estimating policy effects on the same data would lead to overfitting bias, we also provide a data-adaptive estimation framework, which allows for both threshold discovery and policy estimation. Simulation studies show asymptotic convergence to the optimal exposure region and to the true effect of an intervention. We demonstrate how our method identifies true interactions in a public synthetic mixture data set. Finally, we applied our method to NHANES data to discover metal exposures that have the most harmful effects on telomere length. We provide an implementation in the CVtreeMLE R package.
@@ 1 Introduction
Ideally, the analyst would have the full multidimensional dose-response relationship, but this is complex due to unrealistic assumptions and nonpathwise differentiability, complicating robust confidence interval construction. To address this, we suggest categorizing joint exposures to assess the impact of specific exposure thresholds on average outcomes. If our aim is to quantify the effect of imposing certain exposure thresholds, a sensible approach is to categorize the joint exposure and compare the average outcome if everyone were forced to have exposures within a region compared to the observed average outcome under observed exposures. Because this simulated policy is based on a region discovered in the exposure space, we call our target parameter the average regional exposure effect (ARE). This approach is helpful because interpretable categories can be defined, such as $(A_{1}<a_{1})\&(A_{2}<a_{2})$, where $a_{i}$ are specific values in $A$ that are of clear interest to policy makers. Identification assumptions are also more transparent in this setting. In this context, we seek to categorize the exposure space and find the regions with the expected minimum (or maximum) population outcome. However, this intervention that effectively forces people to a region is ambiguous. The post-intervention probability density of exposure $A$ is concentrated in the given region, but there are infinitely many different policies that achieve this. Our choice preserves relative natural selection preferences within the region, per covariate stratum, which is a natural option and lends itself to easy estimation with standard tools. However, the thresholding of the exposure space that minimizes/maximizes the population average outcome is not known a priori. Data must be used to determine which regions are relevant. To estimate this oracle region of the exposure space, we develop a recursive partitioning algorithm paired with g-computation to greedily search the exposure space. Once the thresholds have been identified, we obtain the most precise inference possible for the effects using a doubly robust and efficient estimator, targeted maximum likelihood estimation (TMLE).
@@
Using the same data to define thresholds and estimate effects in the outcome minimum region can lead to biased results. We address this problem by splitting the data, estimating thresholds on one part, and estimating counterfactual effects (given the fixed thresholds) on the other. To efficiently use all the data, we repeat this splitting process in a round-robin fashion (K-fold cross-validation). We build this framework on previous work related to data-adaptive parameters Hubbard et al. [2016] and cross-validated targeted minimum loss-based estimation (CV-TMLE) Zheng and van der Laan [2010], our method, called CVtreeMLE, is a novel approach for estimating the joint impact of a mixed exposure by using CV-TMLE that guarantees consistency, efficiency, and multiple robustness despite using highly flexible learners (ensemble machine learning) to estimate a data-adaptive parameter. A crucial challenge in this process is the inconsistency of the data-adaptive parameter. This is especially evident in the case of variable importance methods, which use quantiles of variables and estimates variable importance based on the average treatment effect (ATE), measuring quantiles with the highest expected mean against the lowest. Hubbard et al. [2018] There is potential variability of the parameters identified (quantiles) across different data subsets that are identified as the maximum or minimum in the respective fold. Such variability can complicate the interpretation of these data-adaptive target parameters. Our current work seeks to address this inherent challenge of a mixture of possibly different regions found across the folds for our target parameter. We not only provide pooled estimates common to data-adaptive target parameters for the maximizing/minimizing W-controlled exposure region, which may be an amalgamation of different regions due to variability in findings across the folds, but also introduce a new approach to pool estimates across different exposure sets. For example, if thresholds are found for $(A_{1}<a_{1})\&(A_{2}<a_{2})$ in 6 of 10 folds, because thresholds are found for this exposure set $A_{1}$ and $A_{2}$, we pool these estimates for this exposure set.
@@
This manuscript is structured as follows: Section 2 discusses semiparametric methodology and redefines the population intervention effect as the average regional effect (ARE). Section 3 focuses on ARE estimation and inference, while Section 4 introduces the oracle target parameter. Section 5 delves into cross-estimation, including pooling methods. Simulation studies demonstrating the estimator’s properties are in Section 6. Section 7 applies CVtreeMLE to NIEHS data and compares it with other methods, section 7.2 we apply our method to NHANES data. The final sections (8 and 9) detail the CVtreeMLE software and conclude with a discussion.
@@ 2.2 Defining the Differential Effect Given Regional Exposure
$=E[E[Y|\mathcal{A}=1,W=w]]$
What we have shown is that this parameter is a population average of some function $Q(\mathcal{A},W)=E[Y|\mathcal{A},W]$ when $\mathcal{A}$ is forced to 1. For any value of $w$, $Q(1,w)$ is a particular convexly weighted average of the causal dose-response curve across the different exposure levels, thus collapsing it down to a single number for each $w$.We want to compare the average outcome if all individuals had exposures within the region $\mathcal{A}$ with the average observed outcome under the observed exposure distribution. This defines our parameter of interest, which we call the average regional effect (ARE): $\psi=E[Q(\mathcal{A}=1,W)]-E[Y]$, representing the difference in the average outcomes if we forced exposure to self-selection within $\mathcal{A}$ versus the observed outcomes. In most applications, it is not known a priori which region $\mathcal{A}$ should be set. For example, we may not know how various chemicals or drugs interact and how to set safe limits for all of them. Therefore, $\mathcal{A}$ itself is in practice something that should be estimated to maximize some objective. That said, for the purposes of establishing our theory, we should first imagine $\mathcal{A}$ as known and fixed. We will then show how we can choose what policy to enact (i.e., choose $\mathcal{A}$) while also unbiasedly estimating its effect. We first discuss the assumption necessary for the statistical quantity estimable from sample data to have causal interpretations, then we move on to discuss how $\mathcal{A}$ can be determined.
2.3 Identification and Causal Assumptions
Our target parameter is defined in terms of the observable distribution and does not represent a causal quantity without additional assumptions. To add a causal interpretation, we define the causal ARE as $\psi^{*}=E[Y(A)|\mathcal{A}=1]-E[Y]$. Standard conditioning arguments show $\psi^{*}=\psi$ when the conditional randomization assumption, $A\perp Y(a)\mid W$ for all $a\in\mathcal{A}$ and positivity: $P(\mathcal{A}=1|W)>0$ for all $w$, are met. The randomization assumption, asserts that the potential outcomes are independent of exposure within covariate-defined strata, indicating that there are no unmeasured confounders. The positivity assumption, $P(\mathcal{A}=1|W)>0$ for all $w$, ensures that all exposure levels and covariate strata are represented in the data, critical for accurate effect estimation. By satisfying these conditions, we can estimate the causal ARE reliably, focusing on obtaining the observable ARE efficiently without additional assumptions such as linearity. Although our identification assumptions may not always hold in all applications, we can at least eliminate model misspecification bias and minimize random variation. Once we have established how to estimate the ARE for a fixed region, we will turn our attention to the problem of finding a good region $\mathcal{A}$ and finally how to do that without incurring selection bias in estimating the ARE for that region.
3 Estimating ARE with TMLE
In the previous sections, we established that the causal ARE is equivalent to the observable ARE $E[E[Y|\mathcal{A}=1,W]-E[Y]]$ under standard identification assumptions. Therefore, to estimate it, all we need to do is 1) create a new binary random variable $\mathcal{A}_{i}=1_{\mathcal{A}}(A_{i})$ and 2) proceed as if we were estimating the observable average counterfactual effect $E[E[Y|\mathcal{A}=1,W]]$ from the observational data structure $(Y,\mathcal{A},W)$ where $\mathcal{A}$ is a simple binary “exposure”.
We can adapt techniques for estimating the ATE $E[E[Y|\mathcal{A}=1,W]]-E[E[Y|\mathcal{A}=0,W]]$ to construct ARE estimators that offer provable unbiasedness (assuming that there is no bias from potential violations of identification assumptions) and efficiency. These estimators are optimized for minimal sampling variance and possess the "doubly robust" property, as highlighted in Zheng and van der Laan [2010], Zivich and Breskin [2021]. Both Augmented Inverse Propensity-Weighting (AIPW) and Targeted Maximum Likelihood Estimation (TMLE) stand out as assumption-free and data-efficient methods when used with machine learning and cross-fitting. Although they often produce similar results, TMLE has been found to perform more efficiently with smaller samples Luque-Fernandez et al. [2018], Smith et al. [2022], van der Laan and Rose [2011, 2018], making it a generally preferable approach. For this reason, we construct a TMLE estimator for the ARE. The TMLE estimator is inspired by the fact that if we knew the true conditional mean $Q(\mathcal{A},W)=E[Y|\mathcal{A},W]$, we could estimate the ARE with the empirical average $\frac{1}{n}\sum_{i}Q(1,W_{i})-Y_{i}$. Of course, we do not know $Q$, but we can estimate it by regressing the outcome $Y$ on the exposure $\mathcal{A}$ and the covariates $W$. However, a detailed mathematical analysis shows that we incur bias if we use our estimate $\hat{Q}$ instead of the truth. This bias might decrease as the sample size increases, but it dominates relative to random variability, making it impossible to establish p-values or confidence intervals. TMLE solves this problem by computing a correction to the regression model $\hat{Q}$ that removes this bias. In other words, “targets” the estimate $\hat{Q}$ to the parameter of interest (here, ARE). We use targeted learning to de-bias our initial predictions of the outcome if all individuals were exposed to exposure levels within the region. The process is as follows: 1. Use cross-validated ensembles of machine learning algorithms (a “super learner”) to generate estimates of the conditional means of treatment: $\hat{g}(\mathcal{A}=a,W)\approx P(\mathcal{A}=a|W)$ (i.e. propensity score) and outcome: $\hat{Q}(\mathcal{A},W)\approx E[Y|\mathcal{A},W]$. 2. Regress $Y$ (scaled to $[0,1]$) onto the “clever covariate” $H_{i}=\frac{\pi(\mathcal{A}_{i})}{\hat{g}(1,W_{i})}$ using a logistic regression with weights $\text{logit}(\hat{Q}(\mathcal{A},W))$. Here, $\pi(\mathcal{A}_{i})$ is the marginal probability of being in the region. The (rescaled) output of this is our target regression model $\hat{Q}^{*}$. That is, using $\epsilon$ from this least favorable model, we debias our predictions for $\hat{Q}(\mathcal{A}=1,W)$. 3. Compute the plug-in estimate using the targeted model: $\hat{\psi}=\frac{1}{n}\sum_{i}\hat{Q}^{*}(1,W_{i})-Y_{i}$. An estimated standard error for $\hat{\psi}$ is given by: $\hat{\sigma}^{2}=\frac{1}{n^{2}}\sum_{i}\left[\left(\frac{\pi(\mathcal{A}_{i})}{\hat{g}(1,W_{i})}\right)\bigg{(}Y_{i}-\hat{Q}^{*}(1,W_{i})\bigg{)}+\bigg{(}\hat{Q}^{*}(1,W_{i})-Y_{i}-\hat{\psi}\bigg{)}-\bigg{(}Y_{i}-E[Y]\bigg{)}\right]^{2}$.
With corresponding 95% confidence interval $\hat{\psi}\pm 1.96\hat{\sigma}$. Here, two terms before the minus are the efficient influence function (EIF) for the parameter $E[E[Y|\mathcal{A}=1,W]]$ and the third term is the EIF for the mean of the random variable $Y$. Explaining why the targeting step takes the form of a logistic regression and how the estimated standard error is derived is beyond the scope of this work. van der Laan and Rubin [2006], van der Laan and Rose [2011, 2018] offer explanations targeted to audiences with varying levels of mathematical sophistication. To obtain these estimates, we need only to specify the ensemble of machine learning algorithms used to estimate the propensity and initial outcome regressions $\hat{g}$ and $\hat{Q}$. Theoretical guarantees hold as long as a sufficiently rich library is chosen. To estimate the ARE, we must also specify the region $\mathcal{A}$ so that we can compute our binary “exposure” variable. The issue of course is that we have been treating $\mathcal{A}$ as a known region, whereas in many applications the important question is determining what guidelines to impose in the first place. This is the focus of the next section.
4 Finding a Good Exposure Region
@@
4.1 G-Computation Decision Tree

### v3

@@ Abstract
Traditional regulations of chemical exposure tend to focus on single exposures, overlooking the potential amplified toxicity due to multiple concurrent exposures. We are interested in understanding the average outcome if exposures were limited to fall under a multivariate threshold. Because threshold levels are often unknown a priori, we provide an algorithm that finds exposure threshold levels where the expected outcome is maximized or minimized. Because both identifying thresholds and estimating policy effects on the same data would lead to overfitting bias, we also provide a data-adaptive estimation framework, which allows for both threshold discovery and policy estimation. Simulation studies show asymptotic convergence to the optimal exposure region and to the true effect of an intervention. We demonstrate how our method identifies true interactions in a public synthetic mixture data set. Finally, we applied our method to NHANES data to discover metal exposures that have the most harmful effects on telomere length. We provide an implementation in the open-source CVtreeMLE R package.
@@ 1 Introduction
Ideally, the analyst would have the full multidimensional dose-response relationship, but this is complex due to unrealistic assumptions and nonpathwise differentiability, complicating robust confidence interval construction. To address this, we suggest categorizing joint exposures to assess the impact of specific exposure thresholds on average outcomes. If our aim is to quantify the effect of imposing certain exposure thresholds, a sensible approach is to categorize the joint exposure and compare the average outcome if everyone were forced to have exposures within a region compared to the observed average outcome under observed exposures. Because this simulated policy is based on a region discovered in the exposure space, we call our target parameter the attributable regional exposure effect (ARE). This approach is helpful because interpretable categories can be defined, such as $(A_{1}<a_{1})\&(A_{2}<a_{2})$, where $a_{i}$ are specific values in $A$ that are of clear interest to policy makers. Identification assumptions are also more transparent in this setting. In this context, we seek to categorize the exposure space and find the regions with the expected minimum (or maximum) population outcome. However, this intervention that effectively forces people to a region is ambiguous. The post-intervention probability density of exposure $A$ is concentrated in the given region, but there are infinitely many different policies that achieve this. Our choice preserves relative natural selection preferences within the region, per covariate stratum, which is a natural option and lends itself to easy estimation with standard tools. However, the thresholding of the exposure space that minimizes/maximizes the population average outcome is not known a priori. Data must be used to determine which regions are relevant. To estimate this oracle region of the exposure space, we develop a recursive partitioning algorithm paired with targeted learning to greedily search the exposure space using a fast machine learning algorithm for propensity score and outcome regression. Once the thresholds have been identified, we obtain more precise inference for the effects using a cross-validated targeted maximum likelihood estimation (CV-TMLE), with a larger set of algorithms in the ensembles used to estimate the propensity and outcome regression.
@@
Using the same data to define thresholds and estimate effects in the outcome minimum region can lead to biased results. We address this problem by splitting the data, estimating thresholds on one part, and estimating counterfactual effects (given the fixed thresholds) on the other. To efficiently use all the data, we repeat this splitting process in a round-robin fashion (K-fold cross-validation). We build this framework on previous work related to data-adaptive parameters Hubbard et al. [2016] and cross-validated targeted minimum loss-based estimation (CV-TMLE) Zheng and van der Laan [2010], our method, called CVtreeMLE, is a novel approach for estimating the joint impact of a mixed exposure by using CV-TMLE that guarantees consistency, efficiency, and multiple robustness despite using highly flexible learners (ensemble machine learning) to estimate a data-adaptive parameter. A crucial challenge in this process is the inconsistency of the data-adaptive parameter. If the region found to minimize/maximize the expected outcome is not the same across all folds the interpretability of the pooled parameter is difficult, although fold specific results still allow for inference Hubbard et al. [2018].
@@
This manuscript is structured as follows: Section 2 discusses semiparametric methodology and redefines the population intervention effect as the attributable regional effect (ARE). Section 3 focuses on ARE estimation and inference, while Section 4 introduces the oracle target parameter. Section 5 delves into cross-estimation, including pooling methods. Simulation studies demonstrating the estimator’s properties are in Section 6. Section 7 applies CVtreeMLE to NIEHS data and compares it with other methods, section 7.2 we apply our method to NHANES data. The final sections (8 and 9) detail the CVtreeMLE software and conclude with a discussion.
@@ 2.2 Defining the Differential Effect Given Regional Exposure
We ask about the differential policy effect of allowing self-selection within a region of the exposure space. Let us consider a situation where there is a community with houses and families have certain probabilities of moving to live in each house. A state enacts a law that restricts housing options to those more than 5 km from a chemical plant. In some situations, such as housing vouchers, it may make sense to say that the new distribution of housing choices under restriction is simply a scaled-up version of the density with no restriction. In this case, the preference profile of houses in the allowed region may change proportionally compared to the unrestricted distributions. This assumption is only realistic in certain circumstances, although it is implicit when we binarize any exposure data. Lee et al. [2024]
For example, if a family were twice as likely to choose House 1 as they were to choose House 2, that relative preference would continue to hold regardless of whether House 3 was an option. But, of course, this may not always be the case and preferences for House 1 and 2 could be different under the House 3 restriction. This is a special kind of modified treatment policy. Formally, let $\mathcal{A}\subset\mathbb{R}^{m}$ be a subset of the exposure space. For example, assume that $\mathcal{A}$ represents allowable neighborhoods or doses of drugs/pollutants that have been deemed safe for combination. Let $T$ denote the binary random variable $1_{\mathcal{A}}(A)$ indicating whether an observation complied with the policy. Let $\pi_{\mathcal{A}}(w)=P(A\in\mathcal{A}\mid W=w)$ be the probability that an observation with covariates $w$ naturally self-selects treatment in the requested region $\mathcal{A}$. We define the modified treatment variable $\tilde{A}$ to represent the exposure distribution once all exposures are forced to self-select within the region $\mathcal{A}$. The modified treatment has the following density:
$p_{\tilde{A}\mid W}(a,w)=\frac{1_{\mathcal{A}}(a)}{\pi_{\mathcal{A}}(w)}p_{A\mid W}(a,w),$
which preserves the relative self-selection preferences for each available exposure and sets the preferences for "outlawed" exposures to zero. Now we can define the expected population outcome if we were to impose this policy:
@@
$=E[E[Y\mid T=1,W=w]].$
What we have shown is that this parameter is a population average of some function $Q(\mathcal{A},W)=E[Y\mid\mathcal{A},W]$ when $\mathcal{A}$ is forced to 1. For any value of $w$, $Q(1,w)$ is a particular convexly weighted average of the causal dose-response curve across the different exposure levels, thus collapsing it down to a single number for each $w$. We want to compare the average outcome if all individuals had exposures within the region $\mathcal{A}$ with the average observed outcome under the observed exposure distribution. This defines our parameter of interest, which we call the average regional effect (ARE): $\psi=E[Q(T=1,W)]-E[Y]$, representing the difference in the average outcomes if we forced exposure to self-selection within $\mathcal{A}$ versus the observed outcomes.
If we do not assume relative self-selection, we would need to account for how each individual’s preference might change under the restriction. For any regulation or restriction, we would need to model the conditional density of exposures in the region given covariates to try and estimate how the exposure density profile would change under restriction. Alternatively, we would need to make assumptions about what the distribution would look like under restriction, for example, perhaps aggregating just below the threshold, leading to higher densities near the boundary and lower densities farther from the threshold. However, we do not know this information a priori. Therefore, in practice, this approach would involve building a density estimator for $p(A\mid W,A\in\mathcal{A})$, which accounts for the restricted set of options for all possible policies, which is computationally infeasible. We would then use these individualized densities to calculate the new exposure levels for each person under the restriction, which would require additional data and assumptions about how preferences change.
For reasons of interpretability, frequency of binarizing exposures in the literature (with aforementioned implicit assumptions), and ease of using existing tools for estimation, we choose to estimate a policy under relative self-selection.
In most applications, it is not known a priori which region $\mathcal{A}$ should be set. For example, we may not know how various chemicals or drugs interact and how to set safe limits for all of them. Therefore, $\mathcal{A}$ itself is, in practice, something that should be estimated to maximize some objective. However, for the purposes of establishing our theory, we should first imagine $\mathcal{A}$ as known and fixed. We will then show how we can choose what policy to enact (i.e., choose $\mathcal{A}$) while also unbiasedly estimating its effect. We first discuss the assumptions necessary for the statistical quantity estimable from the sample data to have causal interpretations, then we move on to discuss how $\mathcal{A}$ can be determined.
2.3 Identification and Causal Assumptions
Our target parameter is defined in terms of the observable distribution and does not represent a causal quantity without additional assumptions. To add a causal interpretation, we define the causal ARE as $\psi^{*}=E[Y(A)|T=1]-E[Y]$. Standard conditioning arguments show $\psi^{*}=\psi$ when the conditional randomization assumption, $A\perp Y(a)\mid W$ for all $a\in\mathcal{A}$ and positivity: $P(T=1|W)>0$ for all $w$, are met. The randomization assumption, asserts that the potential outcomes are independent of exposure within covariate-defined strata, indicating that there are no unmeasured confounders. The positivity assumption, $P(T=1|W)>0$ for all $w$, ensures that all exposure levels and covariate strata are represented in the data, critical for accurate effect estimation. By satisfying these conditions, we can estimate the causal ARE reliably, focusing on obtaining the observable ARE efficiently without additional assumptions such as linearity. Although our identification assumptions may not always hold in all applications, we can at least eliminate model misspecification bias and minimize random variation. Once we have established how to estimate the ARE for a fixed region, we will turn our attention to the problem of finding a good region $\mathcal{A}$ and finally how to do that without incurring selection bias in estimating the ARE for that region.
3 Estimating ARE with TMLE
In the previous sections, we established that the causal ARE is equivalent to the observable ARE $E[E[Y|T=1,W]-E[Y]]$ under standard identification assumptions. Therefore, to estimate it, we need to: 1) create a new binary random variable $T=1_{\mathcal{A}}(A)$ and 2) proceed as if we were estimating the observable average counterfactual effect $E[E[Y|T=1,W]]$ from the observational data structure $(Y,\mathcal{A},W)$ where $\mathcal{A}$ is a simple binary "exposure".
We can adapt techniques for estimating the ATE $E[E[Y|T=1,W]]-E[E[Y|T=0,W]]$ to construct ARE estimators that offer provable unbiasedness (assuming that there is no bias from potential violations of identification assumptions) and efficiency. These estimators are optimized for minimal sampling variance and possess the "doubly robust" property, as highlighted in Zheng and van der Laan [2010], Zivich and Breskin [2021]. Both Augmented Inverse Propensity-Weighting (AIPW) and Targeted Maximum Likelihood Estimation (TMLE) stand out as assumption-free and data-efficient methods when used with machine learning and cross-fitting. Although they often produce similar results, TMLE has been found to perform more efficiently with smaller samples Luque-Fernandez et al. [2018], Smith et al. [2022], van der Laan and Rose [2011, 2018], making it a generally preferable approach.
A comprehensive analysis by Li et al. Li et al. [2022] using ten different nutrition intervention studies showed that TMLE, especially in its cross-validated form (CV-TMLE), consistently provided robust and efficient estimates across various realistic simulation scenarios. The study concluded that the additional layer of cross-validation helps avoid unintentional over-fitting of nuisance parameter functionals, leading to more reliable inferences compared to other estimators. This reinforces the advantage of TMLE in practical applications, emphasizing its superior performance in maintaining efficiency and robustness in finite samples. Therefore, for estimating our data-adaptively identified region, we use the CV-TMLE approach.
The TMLE estimator is inspired by the fact that if we knew the true conditional mean $Q(\mathcal{A},W)=E[Y|T=1,W]$, we could estimate the ARE with the empirical average $\frac{1}{n}\sum_{i}Q(1,W_{i})-Y_{i}$. Of course, we do not know $Q$, but we can estimate it by regressing the outcome $Y$ on the exposure $\mathcal{A}$ and the covariates $W$. However, using our estimate $\hat{Q}$ instead of the truth incurs bias that might decrease as the sample size increases but dominates relative to random variability, making it impossible to establish p-values or confidence intervals. TMLE solves this problem by computing a correction to the regression model $\hat{Q}$ that removes this bias, "targeting" the estimate $\hat{Q}$ to the parameter of interest (here, ARE).
The process is as follows: 1. Use cross-validated ensembles of machine learning algorithms (a "super learner") to generate estimates of the conditional means of treatment: $\hat{g}(\mathcal{A}=a,W)\approx P(\mathcal{A}=a|W)$ (i.e., propensity score) and outcome: $\hat{Q}(\mathcal{A},W)\approx E[Y|\mathcal{A},W]$. 2. Regress $Y$ (scaled to $[0,1]$) onto the "clever covariate" $H_{i}=\pi(\mathcal{A}_{i})/\hat{g}(1,W_{i})$ using a logistic regression with weights $\text{logit}(\hat{Q}(\mathcal{A},W))$. Here, $\pi(\mathcal{A}_{i})$ is the marginal probability of being in the region. The (rescaled) output of this is our target regression model $\hat{Q}^{*}$. 3. Compute the plug-in estimate using the targeted model: $\hat{\psi}=1/n\sum_{i}\hat{Q}^{*}(1,W_{i})-Y_{i}$.
An estimated standard error for $\hat{\psi}$ is given by:
$\hat{\sigma}^{2}=\frac{1}{n^{2}}\sum_{i}\left[\left(\frac{\pi(\mathcal{A}_{i})}{\hat{g}(1,W_{i})}\right)\left(Y_{i}-\hat{Q}^{*}(1,W_{i})\right)+\left(\hat{Q}^{*}(1,W_{i})-Y_{i}-\hat{\psi}\right)\right]^{2}.$
With the corresponding 95% confidence interval $\hat{\psi}\pm 1.96\hat{\sigma}$.
Explaining why the targeting step takes the form of a logistic regression and how the estimated standard error is derived is beyond the scope of this work. van der Laan and Rubin [2006], van der Laan and Rose [2011, 2018] offer explanations targeted to audiences with varying levels of mathematical sophistication. To obtain these estimates, we need only to specify the ensemble of machine learning algorithms used to estimate the propensity and initial outcome regressions $\hat{g}$ and $\hat{Q}$. Theoretical guarantees hold as long as a sufficiently rich library is chosen. To estimate the ARE, we must also specify the region $\mathcal{A}$ so that we can compute our binary "exposure" variable. This is the focus of the next section.
4 Finding a Good Exposure Region
Our target parameter $\psi$ depends on the region $\mathcal{A}$ to which we restrict the exposures (and, as such, should be denoted $\psi_{\mathcal{A}}$ to be precise). When such a region is not already provided (e.g. there are no existing thresholds for a multivariate exposure), it is natural to want to use the data to find a region $A$ that would optimize population outcomes. In the case where the high value of $Y$ is considered harmful, the oracle parameter we seek can be formally described as
@@ 4 Finding a Good Exposure Region
or equivalently $\underset{\mathcal{A}}{\mathrm{argmin}}\ E[E[Y|T=1,W]]$.
When considering this oracle parameter, we also want to ensure the positivity assumption, specifically that the conditional probability of $\mathcal{A}$ given $W$ (i.e., $P(\mathcal{A}|W)$) is bounded away from 0 and 1, we want to avoid extreme propensity scores that could lead to instability or infinite variances in the estimates. This requirement reflects the need for sufficient variability for being inside and outside the exposure region across levels of covariates to make reliable causal inferences.
The oracle parameter in the given context is described as minimizing the expected value of the outcome $Y$ conditional on being in an exposure region $\mathcal{A}$ and given covariates $W$. To incorporate the boundedness of $P(\mathcal{A}|W)$, we adjust the formulation to explicitly ensure that the selected region $\mathcal{A}$ does not result in probabilities of exposure that are too extreme. In this way, we can rewrite the oracle parameter, incorporating this boundedness directly into the optimization:
@@
subject to the constraints:
$\epsilon\leq\pi_{\mathcal{A}}(w)\leq 1-\epsilon,\quad\forall w$
Here, a reasonable $\epsilon$ may be 0.001, this ensures that no subgroup of $W$ is almost always or almost never exposed. This prevents the model from leveraging rare but extreme treatment effects that are not generalizable or might be based on spurious associations due to model overfitting or data anomalies.
[…]
Thus, some restriction of the admissible regions is required, e.g. restricting $\mathcal{A}$ to be axis-aligned rectangles in the exposure space (this restriction also benefits the interpretability of the oracle region). Additional restrictions are also required to make the problem sensible when there is a point minimum in the conditional mean $\mu$. For example, consider a unidimensional exposure $a\in\mathbb{R}^{1}$ and let $\mu(a,w)=a^{2}$. In this case, the minimizing region is the single point $\mathcal{A}=\{0\}$. This single-point case is covered by our positivity restriction to the oracle parameter. Given restrictions on the region, we create a greedy approximation of the oracle region using recursive partitioning of the exposure space, wherein a partition is only made if the conditional probability of being exposed to a region is bounded away from 0 and 1, this algorithm is described next.
4.1 Targeted Decision Trees